OpenStudy (anonymous):

can someone help me to solve this using the de moivre's formula [(square root of 6 + square root of 2)+(square root of 6 - square root of 2) i ] raised to 7

OpenStudy (anonymous):

Well, let's state it first, if \[\huge z = r\left[\cos \theta + i\sin \theta\right]\]then \[\huge z^p=r^p\left[\cos \ p\theta + i\sin \ p \theta \right]\]

OpenStudy (anonymous):

@PeterPan : I got Z^7 = r^7[ cos 7theta +sin7theta*i. plug into the original one I have r=1 cos 7theta = sqr6 + sqr2 sin 7theta = sqr 6 -sqr2 Am I right?

OpenStudy (anonymous):

Uhh, no :) First thing you want to do is convert the form \[\huge a+bi \rightarrow r\left[\cos \theta + i\sin \theta\right]\] It's almost exactly the same as converting rectangular coordinates to Polar :)

OpenStudy (anonymous):

yes, so a = rcos theta

OpenStudy (anonymous):

This isn't your question, @Hoa I wonder why you take an interest in it :> And no, things are not that simple...

OpenStudy (anonymous):

Well, actually, yes \[a = r\cos \theta\]

OpenStudy (anonymous):

hihi... just long time no see that stuff, want to review.

OpenStudy (anonymous):

True enough \[\huge a = r\cos \theta\]But we're not converting from polar to rectangular, in fact, the other way around :P

OpenStudy (anonymous):

the whole thing power to the 7

OpenStudy (anonymous):

First... convert to polar form. Let's get r first, it's easier :P \[\huge r = \sqrt{a^2 + b^2}\]Where \[\huge a = \sqrt6 + \sqrt 2\] \[\huge b = \sqrt6 - \sqrt2\]

OpenStudy (anonymous):

So clearly, \[\huge a^2 + b^2 = 16\]

OpenStudy (anonymous):

ok, how about raise to 7

OpenStudy (anonymous):

Just a sec... It implies that r = 4 Let's leave that be. What's theta?

OpenStudy (anonymous):

actually, i lost!!! just follow you

OpenStudy (anonymous):

Well, experience might tell you that since a and b are both positive, that \[\huge 0 \le \theta \le \frac{\pi}{2}\]

OpenStudy (anonymous):

my answer is 4 raised to 7 cis 7pi/6